OpenStudy (kmallette09):

The square of a number is equal to 10 less than 7 times that number. What are the two possible solutions? Which of the following equations is used in the process of solving this problem? x 2 - 7x + 70 = 0 x 2 - 7x + 10 = 0

OpenStudy (kmallette09):

please help

OpenStudy (kmallette09):

dont just give the answer

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (kmallette09):

not really hope u could help

OpenStudy (skullpatrol):

The square of a number is equal to 10 less than 7 times that number. Let x = a number $$\Huge x^2 = 7x - 10$$ The square of a number is equal to 10 less than 7 times that number

OpenStudy (skullpatrol):

Does that^ make sense?

OpenStudy (kmallette09):

can u explain a little better

OpenStudy (skullpatrol):

10 less than 7 times that number. means 7x — 10

OpenStudy (skullpatrol):

Read it like "10 less, than 7x"

OpenStudy (kmallette09):

so is it is x squared- 7x + 10 = 0

OpenStudy (skullpatrol):

Nope.

OpenStudy (skullpatrol):

The square of a number is equal to 10 less than 7 times that number. Let x = a number $$\Huge x^2 = 7x - 10$$ The square of a number is equal to 10 less than 7 times that number

OpenStudy (kmallette09):

oh ok

OpenStudy (skullpatrol):

Now add 10 to both sides

OpenStudy (skullpatrol):

and subtract 7x from both sides

OpenStudy (kmallette09):

so 7x+10

OpenStudy (skullpatrol):

Start here $$\Huge x^2 = 7x - 10$$

OpenStudy (kmallette09):

so is it x 2 - 7x + 70 = 0

OpenStudy (skullpatrol):

Nope.

OpenStudy (kmallette09):

x 2 + 7x - 10 = 0 so then its this right here

OpenStudy (kmallette09):

right

OpenStudy (skullpatrol):

Yes.

OpenStudy (kmallette09):

thanks

OpenStudy (skullpatrol):

Thanks for trying to learn :-)